# Cocoa: Co-Planning and Co-Execution with AI Agents

### v3

@@ 9. Limitations and Future Work
Based on our formative study, we built Cocoa as a document editor because documents are natural sites of planning for researchers and present an ideal environment for agent interaction (Section 3.2.1). However, the linear nature of a document also has its limitations and does not support “forking” plans and iterating on versions of a plan in parallel. These types of interactions are better supported by node-based canvases, which have been gaining popularity as a means of sensemaking and creative exploration with LLMs (e.g., (Suh et al., 2023; Pu et al., 2024; Arawjo et al., 2023)). Future work can explore new interaction techniques for interactive plans. The default assignment of user and agent steps in Cocoa is also a result of researchers’ preferences from our formative study. Past work explored automated methods for a model to decide when to defer a task to a human expert versus acting on its own (Mozannar and Sontag, 2020). Future work may investigate applications of these methods in interactive plans and/or leverage inference-time scaling to improve plan quality (OpenAI, 2024; Muennighoff et al., 2025).
@@
The underlying agent used in Cocoa also had some technical limitations. Since it was not multimodal, it was not capable of taking visual content (e.g., figures or slides) as input. The agent was also not capable of executing code—a capability which a couple of participants inquired about during the study. While it is sufficient for this current work to explore interactive plans as a novel interaction paradigm, future work can expand their utility with multimodal agents and/or agents that can execute code.
@@
Another limitation is the demographics of our participants. Our participants were researchers in CS and CS-adjacent areas. Research culture and incentives specific to CS may bias our results and limit our imagination of how interactive plans can be used. P10, who used to work in wet labs in their undergraduate research, shared that a useful application of interactive plans would be helping wet lab researchers walk through the steps necessary to prepare for lab experiments. Future work may investigate novel applications of interactive plans in domains beyond CS.
While our work only explored the use of interactive plans by a single user, one interesting future direction is to explore interactive plans in multi-user collaborative scenarios. Interactive plans can potentially coordinate the efforts between multiple human users and an AI agent. Project documents are often highly collaborative artifacts. Interactive plans can play a similar role to existing task management tools like Trello [^11] https://trello.com/ that serve as coordinative artifacts (Bardram and Bossen, 2005; Schmidt and Wagner, 2002) and broaden common ground (Olson and Olson, 2000) in teams. They can help plan to-dos for the team, consolidate outputs in a central location, and display progress as items in the plan get completed.
Novel challenges arise for the agent in these collaborative scenarios. When proposing plans, the agent not only has to reason about whether a step should be assigned to a human user, but also who that will be. To do this, the agent may leverage context in the form of populated profiles of team members (e.g., researchers’ webpages and publication records, LinkedIn profiles outside of research). The agent may also infer expertise from the edit history of individuals within the document, similar to techniques for socially grounding AI chatbots in group chats using conversation history (Wang et al., 2024b). More generally, challenges in human-agent communication get exacerbated in collaborative settings. For example, the question of “which preferences should the agent respect?” (Bansal et al., 2024) is already challenging in a single-user setting because it requires the agent to elicit, preserve, and reason over task context and user preferences simultaneously. Shifting to collaborative settings necessitates keeping track of multiple sets of contexts and preferences and how they interact with each other. Collaborative use of interactive plans thus leave many open questions for future work.
@@ 10. Conclusion
In this paper, we presented Cocoa, an interactive document editing environment for scientific researchers to fluidly collaborate with an AI agent to tackle open questions and tasks within their research projects. Cocoa introduced a new interaction design pattern—interactive plans—that enabled Co-planning and Co-execution between a researcher and an AI agent.
Our lab ($n=16$) and field deployment ($n=7$) studies showed that participants not only engaged in collaborative planning and execution in Cocoa, but also interleaved co-planning and co-execution to progress on their tasks. Compared to a strong baseline with a more familiar chat interface, Cocoa enhanced agent steerability without sacrificing ease of use. In the lab study, participants’ main mode of co-execution was refining agent outputs, but when Cocoa was integrated into their research workflows in the field over a 7-day period, participants took advantage of user steps to guide the agent with their expertise and exert more of their agency. Overall, our work demonstrates the potential of interactive plans as a novel design pattern to flexibly mediate human and AI agency. Importantly, our work offers many practical takeaways for the design of agentic AI systems, such as leveraging user steps and advanced planning for agent cost management and oversight.
Agentic AI systems have exciting potential to transform our digital experiences and advance long-standing visions held by HCI and AI communities. However, these transformations and advancements also demand renewed attention to strategies for ensuring effective collaboration between human users and AI agents in the real world. Our work advances emerging efforts in this new frontier of human-AI collaboration.
@@ Acknowledgements.
We extend a warm thanks to all participants from our formative, pilot, and user studies. We also thank Raymond Fok and Shannon Shen for helpful discussions.
Appendix A Frequently Asked Questions
A.1. What’s the novelty of Cocoa compared to “deep research” systems?
It’s true that Cocoa and deep research systems can be used for similar tasks (e.g., literature review) and might even obtain similar results. However, deep research systems (at least at the time of writing) use fully automated AI workflows and do not accommodate user steering during the research process. Deep research systems may engage in some back and forth with users before the research process to clarify requests, or allow the user to follow up after the process ends, but does not afford user interaction during the process itself.
Why might this be undesirable? Let’s say you want to write a literature review on whether buttons with icons and labels are more usable than buttons without labels, or labels without icons. This is a real example that OpenAI has on their Deep Research release announcement webpage (OpenAI, 2025a). As an HCI researcher, you know of some reputable venues and authors you can rely on, but the agent used sources you might view as less credible (see sources in Deep Research’s response in the UX Design tab of (OpenAI, 2025a)). If Deep Research offered the affordances, you could guide the agent with your expertise during the research process to steer it towards using your trusted sources, which can also help improve the quality of its output. Without these affordances, you could follow up via prompting afterwards, but 1) there’s no guarantee that the agent will reliably follow those new instructions, and 2) re-running the research process is time- and resource-intensive. That’s why we draw on the computational notebook metaphor in Cocoa: interactive plans allow you to iteratively hone and refine the agent’s trajectory and outputs to bolster steerability, control, efficiency, and quality.
A.2. Can a user technically co-plan and co-execute with an agent via chat?
Technically, chat interfaces allow the user to ask the agent for a plan and prompt the agent to regenerate the plan before executing it (which has some aspects of co-planning). The user can also direct the agent to save some steps for themselves and prompt the agent to modify its outputs for steps assigned to the agent (which has some aspects of co-execution). However, these interactions are quite cumbersome in chat and not aligned with users’ mental models of chat interfaces—we did not witness any participants using our baseline system this way. Interactive plans are a fundamentally distinct design pattern designed to support interleaved co-planning and co-execution, and elicits different user behaviors than chat interfaces (see Figures 12 and 13 for evidence of this).
A.3. Do you see Cocoa’s lack of improvement over the baseline for general utility as a negative result?
No, we do not. General utility is a combination of three measures we obtained by asking participants the following Likert scale questions: 1) I’ve developed a better understanding of potential solutions to this particular problem, 2) I obtained new and useful insights from using system, and 3) The system’s outputs were useful for exploring the problem. It is unsurprising that there was no significant difference in answers to these questions, because these aspects are often bottlenecked by the performance of the underlying agent, and we used the same agent in both Cocoa and the baseline. However, we can improve aspects such as steerability with interface-level advances, which we have shown in this work. If general utility had gone up in Cocoa, it would have been a pleasant surprise, but it is understandable if it did not.
A.4. Why is it important to get the user involved? Can’t we just let AI automate everything?
While existing efforts to make agents operate as autonomously as possible without seeking user input may be a compelling research agenda, this approach is not always practical. First, there some tasks humans can more effectively complete than AI, especially if the task requires context or domain expertise the LLM does not have access to. Second, there are some tasks humans may prefer to perform rather than having AI automate it (e.g., if doing the task allows the user to gain new insights or knowledge that they may not obtain otherwise). Third, having an agent operate for a long period of time without adequate steering mechanisms for users can both be resource-inefficient from a cost perspective and unsettling from a user experience perspective.
The list can go on, but the point here is that agents will be working with and operated by humans in the real world, so we need to think more deeply about how humans and agents can collaborate effectively. Our work is one product of such thinking.
A.5. Why didn’t you use ChatGPT or a similar publicly available AI chatbot for your baseline?
Our baseline’s interface closely matches that of ChatGPT and other AI chatbots, but the underlying system is different. We use an LLM agent powered by GPT-4o, but specifically trained it to use the Semantic Scholar API as tools to reliably perform research-related tasks such as searches for papers/authors/topics, sorting papers by various metrics including citation count, summarizing papers, comparing papers, retrieving prominent papers by a particular author, etc. See Section 4.4 for technical details. However, our main contribution is not building a new LLM agent nor advancing agent capabilities—rather, it is to explore cost and benefit tradeoffs between interaction techniques for humans to work with AI agents. We use the same underlying agent in Cocoa, so we kept this factor constant across our system and baseline.
A.6. I’m a developer who builds agentic AI applications. What are some practical takeaways your work has for me?
Our Discussion and Future Work section is where you’d find most of our practical takeaways, so we’d encourage you to read that. In addition, here are two high-level takeaways to consider:
  • Users may not want agents to automate every step of workflows for many reasons, whether it’s to have more ownership over their work, or feeling that the agent is incompetent at performing a particular task. Strategic and configurable automation is key to building great user experiences for agents.
  • Agents can be prohibitively expensive to run, and this may be one of the major bottlenecks for widespread adoption. Because agents will inevitably work with humans in real-world settings, it’s important to identify opportunities for human input to lower cost AND improve system performance.
A.7. Can I use Cocoa and/or the baseline?
We are continuing to iterate on Cocoa and the baseline, so we are not yet planning a public release. However, we encourage you to check out Ai2’s ScholarQA [^12] https://scholarqa.allen.ai/chat and PaperFinder [^13] https://paperfinder.allen.ai/chat tools, which share some similar features.
Appendix B Formative Study Participants
@@
Appendix C Cocoa Formative Study Procedure Details
@@
C.1. Project document activity
@@ C.2. Probe activity
C.3. Concluding interview
The study ended with a concluding interview, where we asked participants about their experiences, perceptions, and desiderata for the AI assistant, along with preferred ways to interact with it. Each participant was given a \$35 USD honorarium after the study. The study was reviewed and approved by our organization’s internal IRB.
Appendix D Formative Study Data Analysis
@@ Appendix H Lab Study Self-Evaluation Form Questions
Appendix I Prompts
I.1. Plan generation
# Instructions
You are a helpful research assistant. You will be given a high-level request by a researcher. Your task is to create a short plan to accomplish that request as effectively and efficiently as possible. The high-level request may already come with a partially-completed plan; if that is the case, help complete the rest of the plan in a coherent and sensible manner. You will create the plan by assembling a series of plan steps. These plan steps can be executed by an AI agent, or by a user, as the user should still be involved when making key decisions. In general, users perform steps that involve higher-level reasoning and synthesis.
Below is a catalogue of plan steps you might use to assemble plans. Variables that are in [square brackets] can be populated with what you think is appropriate and useful. Each plan step must contain the following components:
- **description**: a string that provides a high-level description of the step, to be displayed in the UI
- **actor_user**: a boolean value indicating whether the step should be executed by the user
- **output_format**: a string specifying the format of the output data generated by this step. One of: paper_list, author_list, topic_list, entity_list, text
- **score**: a numerical value with one decimal place from -1.0 to 1.0. A score of 1.0 indicates that the step is to be carried out by a human user, while a score of -1.0 indicates that the step is to be carried out by an AI agent.
## Data types:
Here are the data types you can use for output_format, followed by a brief description of each:
- **paper_list**: a list of research papers, represented by their corpus IDs and other accompanying metadata
- **author_list**: a list of authors, represented by their author IDs and other accompanying metadata
- **topic_list**: a list of topics, represented by their topic IDs and other accompanying metadata. Note that these are different from papers. A paper contains a title, abstract, authors, etc., while a topic is a high-level concept or field of study that may describe a set of papers.
- **entity_list**: a list of entities in natural language, such as concepts, ideas, terms, text snippets, or questions
- **text**: a block of freeform natural language text
## Example user steps (note that you do not have to follow these---these are just some possibilities):
Read relevant papers and note down key insights
- **description**: "Read relevant papers and note down key insights"
- **actor_user**: True
- **output_format**: text
- **score**: 1.0
Iterate approach based on feedback, and jot down any notes
- **description**: "Iterate approach based on feedback, and jot down any notes"
- **actor_user**: True
- **output_format**: text
- **score**: 1.0
Brainstorm cases when this [approach] may fall short for [problem]
- **description**: "Reason about cases when this [approach] is not suitable for [problem]"
- **actor_user**: True
- **output_format**: text
- **score**: 1.0
Reason about approaches to adapt an existing approach to [new context]
- **description**: "Reason about approaches to adapt an existing approach to [new context]"
- **actor_user**: True
- **output_format**: text
- **score**: 1.0
Identify papers that can seed exploration in [area]
- **description**: "Identify papers that can seed exploration in [area]"
- **actor_user**: True
- **output_format**: paper_list
- **score**: 1.0
Run an experiment to check the feasibility of [approach or idea], and jot down any notes
- **description**: "Run an experiment to check the feasibility of [approach or idea], and jot down any notes"
- **actor_user**: True
- **output_format**: text
- **score**: 1.0
Write down desired key contributions
- **description**: "Write down desired key contributions"
- **actor_user**: True
- **output_format**: text
- **score**: 1.0
Reason about how [insights] can be formulated into testable hypotheses, and jot down any notes
- **description**: "Reason about cases when this [approach] is not suitable for [problem]"
- **actor_user**: True
- **output_format**: text
- **score**: 1.0
Iterate on [ideas] based on feedback, and jot down any notes
- **description**: "Iterate [ideas] based on feedback, and jot down any notes"
- **actor_user**: True
- **output_format**: text
- **score**: 1.0
## Example agent steps:
Search for papers that discuss [query] and sort by [criteria]
- **description**: "Search for papers that discuss [query] and sort by [criteria]"
- **actor_user**: False
- **output_format**: paper_list
- **score**: -1.0
Find papers related to [entity]
- **description**: "Find papers related to [entity]"
- **actor_user**: False
- **output_format**: paper_list
- **score**: -1.0
Brainstorm search queries for finding papers relevant to [topic]
- **description**: "Generate search queries for finding papers relevant to [topic]"
- **actor_user**: False
- **output_format**: entity_list
- **score**: -1.0
Answer [question] with information from relevant papers
- **description**: "Answer [question] with relevant papers"
- **actor_user**: False
- **output_format**: text
- **score**: -1.0
Identify authors who have published on [topic]
- **description**: "Identify authors who have published on [topic]"
- **actor_user**: False
- **output_format**: author_list
- **score**: -1.0
Find notable papers written by [author] on [topic]
- **description**: "Find papers written by [author] on [topic]"
- **actor_user**: False
- **output_format**: paper_list
- **score**: -1.0
Suggest some common themes in relevant papers on [topic]
- **description**: "Suggest some common themes between relevant papers on [topic]"
- **actor_user**: False
- **output_format**: text
- **score**: -1.0
Summarize key insights collected thus far
- **description**: "Summarize key insights collected thus far"
- **actor_user**: False
- **output_format**: text
- **score**: -1.0
Provide constructive feedback on [topic], grounded in existing literature
- **description**: "Provide constructive feedback on [topic], grounded in existing literature"
- **actor_user**: False
- **output_format**: text
- **score**: -1.0
Suggest some connections between [topic] and [topic], grounded in existing literature
- **description**: "Suggest some connections between [topic] and [topic], grounded in existing literature"
- **actor_user**: False
- **output_format**: text
- **score**: -1.0
## Examples
Below are some examples of plans created using these plan steps. Note that these examples all use the description field of the steps and user steps (where actor_user is True) have a [user step] label. These are only for guidance and do not have to be strictly followed.
What works are there on tool selection by LLM agents?
- Brainstorm search queries for finding papers relevant to tool selection by LLM agents
- Search for papers using selected search queries and sort by relevance
- [user step] Read relevant papers and note down key insights
- [user step] Write down desired key contributions
- Provide constructive feedback on the key contributions, grounded in existing literature
What datasets exist for fine-tuning LLM agents?
- [user step] Identify papers that seed exploration in fine-tuning LLM agents
- Find papers related to seed papers and sort by relevance
- Answer "what datasets are used for fine-tuning LLM agents?" with relevant papers
- [user step] Brainstorm cases when this dataset may fall short for fine-tuning LLM agents
What are techniques from cognitive science that can inform LLM-based agent architectures?
- Brainstorm a list of topics or concepts relevant to "agent architectures" from cognitive science
- Suggest some connections between these concepts and LLMs, grounded in existing cognitive science literature
- [user step] Reason about how these connections can be formulated into testable hypotheses, and jot down any notes
- Provide constructive feedback on the proposed hypotheses, grounded in existing literature
- [user step] Iterate on hypotheses based on feedback, and jot down any notes
Critique the following idea, using perspectives from prior work: human-in-the-loop LLM agents that enable interactive co-planning between a human user and an LLM agent
- [user step] Identify papers that seed exploration in fine-tuning LLM agents
- Answer "what are some critiques of human-in-the-loop?" with relevant papers
- Summarize key insights collected thus far
- [user step] Synthesize critiques and identify salient research directions to address them
- Provide constructive feedback on the identified research directions, grounded in existing literature
How to perform human evaluation on LLM agent outputs?
- Identify authors who have published on LLM agents and/or human evaluation
- Find notable papers written by relevant authors on the topic and sort by citation count
- Answer "how is human evaluation performed on LLM outputs?" with relevant papers
- [user step] Reason about how to adapt an existing approach to LLM agents, and jot down any notes
- Check for papers closely related to the adapted approach
- [user step] Run an experiment to check the feasibility of the adapted approach, and jot down any notes
## Output formatting (follow closely!)
You will output an array of valid JSON objects that represent a plan, given a request. The fields in the JSON and their contents should match the components of a plan step specified earlier (description, actor_user, output_format, score). You should take the following approach:
1. Assemble the plan with just the descriptions of each plan item, filling in any [square brackets] with the appropriate information
2. Process the plan into the format of a JSON array by converting each description into a JSON object with the corresponding fields
3. You can be creative and generate plan steps with a description outside of the ones provided, especially for user steps. If that is the case, make sure the same requirements for other components of the plan step still apply. For example, assign a reasonable data type to the step
Make sure to keep plans as short, simple, and straightforward as possible, ideally shorter than 5 steps. Once again, be creative with the plans and use the provided context if its relevant. Not all plans will involve searching for papers, so only conduct a paper search if appropriate. Also make sure to return valid JSON. The format of the final output must just be an array of valid JSON objects, no additional text.
I.2. Plan description to agent instruction
You are a helpful planning assistant. You will be provided with 1) contexts that are represented in a structured JSON format, and 2) a description of the current plan step. Each context entry is a previously completed plan step; the entries can be found as objects in an array in the "contexts" field of the JSON. Each context JSON object in the array contains a description of the corresponding plan step, the output of the step, and the steps ID. The most important piece of information in this object is the output. There is also a high-level user request in the user_request field of the top-level JSON, but you should not pay much attention to it unless the unless the context field is an empty array.
Based on the provided information, you will generate a list of requests to an AI assistant that has access to scholarly knowledge and literature. Do not actually complete these requests yourself, they are requests for another AI assistant. For example, when asked to brainstorm search queries, dont actually return search queries yourself. Just repeat the instructions back and supplement with information from context if needed. To come up with these requests, you should follow these steps:
1. **Determine which pieces of context are relevant to the description**: given the description, determine which context entries (specifically which outputs) are relevant to the plan step. You should give more importance to more recent context entries. Context entries towards the end of the list of context entries are more recent. If the description of the current plan step contains mentions of any specific information entities (papers, search queries, concepts, etc.), be sure to take those from the outputs of the most recent entry in which they are available.
2. **Determine how many requests are needed**: based on the description of the current plan step and its relevant context entries, determine how many requests are needed to complete the plan step. Does the description indicate a need to loop over multiple information entities provided in the context? If so, more than one request is needed. If no looping is required, only one request is needed.
3. **Generate a list of requests**: based on how many requests are needed, generate a list of requests, where each request is a string. Each request should be clear, specific, and concise. Here are some more specific instructions to follow:
- Each request should be based off the description initially provided and modified by relevant context identified in step 1.
- Be sure to include relevant paper IDs (for papers) in the request, so the agent can properly retrieve paper information. However, do not include IDs for topics, since the AI assistant cannot retrieve information based on topic IDs.
- The list of requests should be as short as possible. Do not generate any unnecessary requests, and try to complete the plan item with as few requests as possible (e.g., listing many paper IDs for summarization at once). However, be careful with search queries and answering a question from a paper---do not try and combine multiple search queries into one request, or answer a question. Each search query should be a separate and distinct action.
- When being asked to search for a paper, do not include too many search queries, or else it will be too difficult to find papers. Make sure the search query is short and contains no more than 3 key search queries or topics.
- Make sure the request appropriately reflects the description. Do not ask for papers if the description is about a list of topic or concepts.
- For descriptions that ask for answering a question from a paper, include "use paper_qa" and relevant paper IDs in the request. Do not mention paper_qa when generating summaries.
4: **Filter the list of requests**: if there are more than 10 requests, filter the list down to the 5-10 most relevant and meaningful requests. If there are fewer than 10 requests, keep all of them.
You will output a list of strings with the final generated requests, in the form ["request1", "request2", ...]. If only one request is needed, the list will contain just one request. You will output this list as the final output with no additional text.
I.3. Agent output reformatting for displaying in UI
You are a helpful planning assistant. You will be provided with 1) contexts that are represented in a structured JSON format, 2) a description of the current plan step, and 3) the output format of the current plan step. Each context entry is a previously completed plan step; the entries can be found as objects in an array in the "contexts" field of the JSON. Each context JSON object in the array contains a description of the corresponding plan step, the output of the step, and the steps ID. The most important piece of information in this object is the output. There is also a high-level user request in the user_request field of the top-level JSON, but you should not pay much attention to it unless the context field is an empty array.
You will use the current plan step description and relevant context to populate a UI determined by the output format that will allow a user to interactively complete the plan step. To do this, follow these steps:
1. **Determine which pieces of context are relevant to the description**: given the description, determine which context entries (specifically which outputs) are relevant to the plan step. You should give more importance to more recent context entries. Context entries towards the end of the list of context entries are more recent. If the description of the current plan step contains mentions of any specific information entities (papers, search queries, concepts, etc.), be sure to take those from the outputs of the most recent entry in which they are available.
2. **Process the context based on output format**: using relevant context entries, process their outputs by following the formatting instructions specific to each output format:
- **paper_list**: a list of corpusIds from papers. Ex: [’corpusId1’, corpusId2’, ...]
- **author_list**: a list of authorIds. Ex: [’authorId1’, authorId2’, ...]
- **topic_list**: a list of topicIds. Ex: [’topicId1’, topicId2’, ...]
- **entity_list**: a list of entities in natural language. Ex: [’entity1’, entity2’, ...]
- **text**: unstructured, natural language text
If the output format is text, all youll need to do is provide an empty string "" for the user to complete themselves. You will provide the processed outputs from context entries. Only select entries that are highly relevant. Only provide the processed outputs, with no additional text.
Appendix J Deployment Study General Guidelines

### v4

@@ 9. Limitations and Future Work
Based on our formative study, we built Cocoa as a document editor because documents are natural sites of planning for researchers and present an ideal environment for agent interaction (Section 3.2.1). However, the linear nature of a document also has its limitations and does not support “forking” plans and iterating on versions of a plan in parallel. These types of interactions are better supported by node-based canvases, which have been gaining popularity as a means of sensemaking and creative exploration with LLMs (e.g., (Suh et al., 2023; Pu et al., 2024; Arawjo et al., 2023)). Future work can explore new interaction techniques for interleaved co-planning and co-execution. The default assignment of user and agent steps in Cocoa is also a result of researchers’ preferences from our formative study. Past work explored automated methods for a model to decide when to defer a task to a human expert versus acting on its own (Mozannar and Sontag, 2020). Future work may investigate applications of these methods in human-agent systems and/or leverage inference-time scaling (OpenAI, 2024b; Muennighoff et al., 2025) to improve plan quality.
@@
Our baseline mimicked the dominant interaction paradigm in AI-powered research tools: chatbots. While this allowed us to elicit participant feedback on the usability of Cocoa compared to the status quo, we could have designed additional baselines by ablating parts of our design. For example, our baseline could have been an interactive planning interface that did not support stepwise execution. In future work, we can expand our participant pool and run more studies with ablations of Cocoa to pinpoint which particular affordances participants valued most. Further, the underlying agent used in Cocoa also had some technical limitations. Since it was not multimodal, it was not capable of taking visual content (e.g., figures or slides) as input. The agent was also not capable of executing code—a capability which a couple of participants inquired about during the study. Future work can expand Cocoa’s utility with multimodal agents and/or agents that can execute code.
@@
Another limitation is the demographics of our participants. Our participants were researchers in CS and CS-adjacent areas. Research culture and incentives specific to CS may bias our results and limit our imagination of how Cocoa can be used. P10, who used to work in wet labs in their undergraduate research, shared that a promising use case for Cocoa would be helping wet lab researchers walk through the steps necessary to prepare for lab experiments. Future work may investigate novel applications of systems like Cocoa in domains beyond CS.
@@ 10. Conclusion
In this paper, we presented Cocoa, an interactive system for scientific researchers to flexibly collaborate with an AI agent in a document editing environment to tackle open questions and tasks within their research projects. Motivated by a formative study with 9 researchers, Cocoa introduced a new approach for human-agent collaboration—interleaved co-planning and co-execution—by which a human and agent can iteratively create plans of action and execute the plan in a mixed-initiative manner that centers human agency.
Our lab ($n=16$) and field deployment ($n=7$) studies showed that participants benefited from interleaved co-planning and co-execution in the lab and in the field. Compared to a strong baseline with a more familiar chat interface, Cocoa enhanced agent steerability without sacrificing ease of use. In the lab study, participants’ main mode of co-execution was refining agent outputs, but when Cocoa was integrated into their research workflows in the field over a 7-day period, participants took advantage of user steps to guide the agent with their expertise and exert more of their agency. Overall, our work demonstrates the promise of our approach for flexibly mediating human and AI agency. Importantly, our work offers many practical takeaways for the design of human-agent systems, such as leveraging user steps and planning ahead for agent cost management and oversight.
AI agents have exciting potential to transform our digital experiences and advance long-standing visions held by HCI and AI communities. However, these transformations and advancements also demand renewed attention to strategies for ensuring effective and flexible human-agent collaboration. Our work advances efforts in this emerging frontier.
@@ Acknowledgements.
We warmly thank all our participants from our formative, pilot, lab, and deployment studies. We thank our anonymous reviewers for feedback on our manuscript. Finally, we thank Raymond Fok and Shannon Shen for helpful discussions.
Appendix A Formative Study Participants
@@
Appendix B Cocoa Formative Study Procedure Details
@@
B.1. Project document activity
@@ B.2. Probe activity
B.3. Concluding interview
The study ended with a concluding interview, where we asked participants about their experiences, perceptions, and desiderata for the AI assistant, along with preferred ways to interact with it. Each participant was given a \$35 USD honorarium after the study. The study was reviewed and approved by the Allen Institute of AI’s IRB.
Appendix C Formative Study Data Analysis
@@
Appendix D Further Implementation Details
D.1. Underlying LLM Agent
D.1.1. Research support tools
We created a custom tool-calling LLM agent [^13] https://platform.openai.com/docs/guides/function-calling powered by GPT-4o. While GPT-4o provides APIs for managing multi-turn conversation context, the system manages its own context using a JSON schema we developed for storing key insights from papers and other forms of output without storing the full artifact. Previous plan steps and their outputs are also represented in the schema to enrich context. To provide high-quality research support, we incorporated two best-in-class research support functionalities into our agent. The first tool allowed the agent to access The Semantic Scholar Open Data Platform (Kinney et al., 2023) via its publically available search APIs. [^14] https://www.semanticscholar.org/product/api This tool provided the agent access to the full text and metadata (such as authors and publication venues) of millions of research papers, created by a state-of-the-art PDF extraction and knowledge graph normalization pipeline, allowing the agent to search for authors and papers based on topics, titles, and rich metadata (Kinney et al., 2023). The second tool is the Ask this Paper [^15] https://www.semanticscholar.org/product feature on Semantic Scholar which generates paper summaries based on the paper’s full text given an aspect of interest. This tool allowed the agent to summarize a relevant paper based on specific aspects of interest within a plan, and also synthesize specific aspects across many papers. For example, the agent can use the first tool to find papers relevant to a topic of interests from a particular author, then, used the second tool to summarized the findings of each papers, and, finally, synthesize the individual summaries into a compare and contrast statement about the set of papers for the user to review. All prompts used are provided in our Supplementary Materials.
D.2. Guidance for plan generation
We use findings from our formative study to guide the agent’s generation of initial plans. In our formative study, we had asked participants to write down brief plans when interacting with our probe (see Appendix B). Based on participants’ preferences of which steps they prefer assigning to an AI vs. keeping for themselves (F3 from our formative study), we organized these steps into agent and user steps. We then wrote examples of how these steps are composed into plans for tackling particular questions, once again drawing from plans participants wrote in the probe activity. These examples were provided for in-context learning in Cocoa’s system prompt for plan generation (see prompts in our Supplementary Materials). Finally, we also enable Cocoa to learn from interactive plans the user has previously created and edited. When the agent is invoked, Cocoa collects the existing plans in the document and adds them to the in-context learning examples on-the-fly. This allows the user’s co-planning efforts to be reused for similar requests elsewhere in the document.
We also enable the agent to learn from previously created plans by extracting existing plans in the document and integrating them into system context.
@@ Appendix H Lab Study Self-Evaluation Form Questions
Appendix I Deployment Study General Guidelines